# Detecting Wildfires on UAVs with Real-time Segmentation Trained by Larger Teacher Models

### v2

@@ Abstract
Early detection of wildfires is essential to prevent large-scale fires resulting in extensive environmental, structural, and societal damage. Uncrewed aerial vehicles (UAVs) can cover large remote areas effectively with quick deployment requiring minimal infrastructure and equipping them with small cameras and computers enables autonomous real-time detection. In remote areas, however, the UAVs are limited to on-board computing for detection due to the lack of high-bandwidth mobile networks. This limits the detection to methods which are light enough for the on-board computer alone. For accurate camera-based localisation, segmentation of the detected smoke is essential but training data for deep learning-based wildfire smoke segmentation is limited. This study shows how small specialised segmentation models can be trained using only bounding box labels, leveraging zero-shot foundation model supervision. The method offers the advantages of needing only fairly easily obtainable bounding box labels and requiring training solely for the smaller student network. The proposed method achieved 63.3% mIoU on a manually annotated and diverse wildfire dataset. The used model can perform in real-time at $\sim$25 fps with a UAV-carried NVIDIA Jetson Orin NX computer while reliably recognising smoke, demonstrated at real-world forest burning events. Code is available at: https://gitlab.com/fgi_nls/public/wildfire-real-time-segmentation
@@ 1 Introduction
The frequency and intensity of extreme wildfires are increasing on Earth, posing tremendous risks to the environment, ecosystems, and societies [12]. Early detection is an essential tool to prevent large-scale disasters [3]. Uncrewed aerial vehicle (UAV)-based detection is a promising technology with the potential to enable rapid deployment of surveys over large areas with minimal infrastructure. In remote areas, however, the UAVs are limited to on-board computing for detection due to the lack of high-bandwidth mobile networks.
@@
In this work, we propose using segmentation models capable of accurate real-time detection of wildfires with purely UAV-on-board computing resources. We leveraged bounding box-guided pseudo-label supervision from large foundation models to train the specialised models as simplified in Figure 1. We introduce two alternative approaches for supervising the final inference model using either a bounding box-guided zero-shot foundation model teacher or a larger model trained using state-of-the-art bounding box supervision methods. The methods were evaluated on a diverse human-annotated test set which showed their benefits compared to direct bounding box supervision methods. Real-world demonstrations and tests on additional data and hardware settings indicated the methods’ real-world applicability. The results suggest that largely used bounding box detection methods could be turned into segmentation at minimal costs.
@@ 2 Related Work
The ideas of UAV-based wildfire monitoring, detection and localisation date back to the early 2000’s [5, 38, 6]. However, at the dates of these studies, the available computational resources and non-learning-based computer vision methods posed serious limitations. More recent studies have applied bounding box detection to the task of wildfire detection [73, 50, 64, 53]. and some of the most used bounding box detection models include R-CNN [19], Faster R-CNN [48], YOLO [47], with its numerous variants, and FPN [31]. Notably, the single-stage object detection networks, including YOLO and its variants focus on the real-time performance of the task and at best can work with very limited computational resources, even ones that a small UAV can carry.
@@
The bounding box detections, however, can not fully capture the shape of the detected object which, in the case of UAV-based wildfire detection, could affect the localisation accuracy by even thousands of meters depending on the flight altitude and camera angle. This is key to our study, as we wish to develop a system capable of accurately estimating the 3D location of the fire onboard the UAV. Segmentation combats this issue by capturing the shape of the detected object with pixel-level accuracy. Recent segmentation models, such as ERFNet [49], BiSeNet [71], and ENet [43] have been proposed specifically for real-time use. Segmentation has also been applied to wildfire detection [51, 72, 61]. However, while the studies have focused on UAV-captured images, their inference times have not been estimated with UAV-carried computers. Also, the studies have mainly focused on flame detection but wildfire smoke is most often visible far further than any flames. The studies on the segmentation of smoke have been largely limited by the lack of public data, especially such, that is annotated at the pixel level.
@@
Publicly available wildfire datasets include the FLAME dataset [52], its synthetic extension FLAME Diffuser [60] and the Corsican dataset [57], all of which focus on flames. The only publicly available datasets with a focus on the resulting smoke that could be found at the date of this study were the Bounding Box Annotated Wildfire Smoke Dataset Versions 1.0 and 2.0 by AI For Mankind and HPWREN [18] and a small dataset by the Croatian Center for Wildfire Research [28] consisting of only 98 images with varying quality pixel-level annotations.
@@
To tackle the lack of pixel-level annotated wildfire smoke data, we explore weak supervision, which could enable training segmentation models using other than dense pixel-level annotations. Weakly supervised segmentation refers to training segmentation models using non-pixel-level labels or far fewer labels than the number of input pixels. This can be realised with image-level labels, scribbles, individual pixels, or bounding boxes, as in the case of this study. The earlier solutions for box-supervised training of semantic and instance segmentation have used iterative mask proposals [13] or a combination of classical (non-learning-based) and learning-based solutions [26]. Many of the recent studies on weakly supervised segmentation have focused on instance segmentation which refers to models which differentiate between object instances without any semantic or object class information. Some of the recent proposals include Mask Auto-Labeler (MAL) [29], BoxTeacher [9], Semantic-aware Instance Mask Generation [30], and BoxSnake [69]. For wildfire detection, weakly supervised methods have been proposed to generate synthetic data [42, 74], to produce flame and smoke segmentation from classification models [1], and to train smoke classification and flame segmentation models [62]. While some of those models have been suggested for UAV-collected images, their capability for real-time sensing with limited hardware has not been studied.
@@
To obtain real-time capable segmentation models, we suggest using knowledge distillation. Knowledge distillation is a way to train effective specialised networks by minimising the difference between the outputs of a larger model, which is assumed capable of better performance and those of a smaller, distilled, model [22]. Knowledge distillation has also been applied to semantic segmentation model training for different applications in various forms with positive results [34, 44, 68, 67]. Knowledge distillation has also been proposed to improve wildfire detection [54, 65] but only in the form of bounding box detection models.
@@
The recent emergence of prompt-guided segmentation [37] and specific models for the task, such as Segment Anything Model (SAM) [27] and High-Quality SAM (HQ-SAM) [25], provides new tools for generating pixel-level pseudo-labels from weak labels. The use of combined foundation model and weak label supervision has been shown successful with image level supervision [70, 55, 7, 8] alongside Class Activation Maps (CAMs) [75] or language understanding from language-image foundation models, such as Grounded-DINO [33] or CLIP (Contrastive Language-Image Pre-training) [45]. Foundation models have been applied to synthetic annotated wildfire flame image generation [60] but distilling the information to real-time models has not been studied nor has the applicability to UAV-based systems in general. However, the use of combined bounding box and foundation model supervision for training use case-specific models has been proposed for planetary geological mapping [23], cell nuclei segmentation [11], and remote sensing [59]. For cell segmentation of histopathological imagery, such a model has been compared to methods that also use SAM at test time [58]. While the studies do not provide generalised results they suggest that such methods could also be applied to train real-time capable wildfire smoke detection models.
@@ 3 Methods
In this study, we address the gaps in the conducted literature by analysing the performance of PIDNet [66] models trained for the segmentation of wildfire smoke using labels obtained from a larger teacher model. We evaluated different teacher models for training the inference models by comparing the resulting accuracy of the final PIDNets. In the main proposed framework, the teacher model is a Segment Anything Model guided by the bounding box labels to generate pseudo-labels for supervising the inference model. We compared this method to an alternative, where the teacher model is a larger binary segmentation model trained using state-of-the-art bounding box supervision methods for segmentation.
We chose PIDNet as the final inference model due to its capability of real-time performance even with limited computational resources. The PIDNet model is a three-branch model where each branch focuses on a separate task: local semantic features, global semantic features, and boundary detection. The predictions are combined for the final inference segmentation output. The training method can be formulated as a knowledge distillation scheme with a teacher and a student model. The training scheme is shown in Figure 3. The data, pseudo-labelling, and inference model training are explained in more detail in the following subsections.
@@ 3.1 Data
where $N$ denotes the total number of samples and $A_{n}$ and $B_{n}$ denote the predicted and the label masks of the individual samples respectively. In addition, the accuracy, precision, recall, and $F_{1}$ scores are reported in the main model comparison and their respective equations are provided in Appendix Eqs. 3, 4, 5 and 6.
@@ 3.3 Inference model training
We used AdamW [36] with a starting learning rate of 1e-3 and weight decay coefficient of 1e-2 to optimise the inference models. The models were trained for 50 epochs and the validation loss and mean intersection over union (mIoU) was computed on each epoch, from which the best-performing model was chosen for evaluation on the test set and real-world inference based on the validation set mIoU. Minibatches of 16 images were used for the training and the models were pretrained on ImageNet [14].
@@
where loss functions $l_{0}$ and $l_{1}$ are binary cross entropy losses from the two alternative segmentation outputs of the PIDNet. $l_{2}$ is a weighted boundary-awareness cross entropy loss [56] and $l_{3}$ is a BAS-Loss that is a binary cross entropy loss computed only at the pixels where the boundary output of the model exceeds the threshold $t$. For the weighting of the losses we use the parameters $\lambda_{0}=0.4$, $\lambda_{1}=20$, $\lambda_{2}=1$, $\lambda_{3}=1$ and $t=0.8$ as set empirically in the study introducing the PIDNet model.
@@ 4 Results
Out of the PIDNet variants, PIDNet-M was found to be the best for the task regarding the test set metrics regardless of the teacher model. However, for practical inference, the PIDNet-S was chosen because of the far superior inference speed on the NVIDIA Jetson Orin NX, shown in Table 3 and the more extensive model speed benchmark with hardware setting variation in Table 14 of the supplementary material. We expect the consistently worse performance of the PIDNet-L model to be caused by overfitting to noisy pseudo-labels, caused by the teacher models poorly capturing the human perception of the smoke in some images. Still, it is interesting how this behaviour was reflected in all teacher variants.
@@ 5 Conclusions
The proposed method also heavily relies on the quality of the SAM-generated pseudo-labels, which as indicated by the results, were not perfect. Thus, it would be of high value to further explore how the masks could be improved. Some simple strategies, such as adding variation to the prompts by random shifting of the manually generated bounding boxes or by assigning randomly chosen points inside the box as prompts could prove beneficial. The variationally generated masks could be used either as label augmentations or to produce more consistent pseudo-labels. Another possibility would be to fine-tune SAM specifically for the smoke segmentation task.

### v3

@@ Abstract
Early detection of wildfires is essential to prevent large-scale fires resulting in extensive environmental, structural, and societal damage. Uncrewed aerial vehicles (UAVs) can cover large remote areas effectively with quick deployment requiring minimal infrastructure and equipping them with small cameras and computers enables autonomous real-time detection. In remote areas, however, detection methods are limited to onboard computation due to the lack of high-bandwidth mobile networks. For accurate camera-based localisation, segmentation of the detected smoke is essential but training data for deep learning-based wildfire smoke segmentation is limited. This study shows how small specialised segmentation models can be trained using only bounding box labels, leveraging zero-shot foundation model supervision. The method offers the advantages of needing only fairly easily obtainable bounding box labels and requiring training solely for the smaller student network. The proposed method achieved 63.3% mIoU on a manually annotated and diverse wildfire dataset. The used model can perform in real-time at $\sim$25 fps with a UAV-carried NVIDIA Jetson Orin NX computer while reliably recognising smoke, as demonstrated at real-world forest burning events. Code is available at: https://gitlab.com/fgi_nls/public/wildfire-real-time-segmentation
@@ 1 Introduction
The frequency and intensity of extreme wildfires are increasing on Earth, posing tremendous risks to the environment, ecosystems, and societies [12]. Early detection is an essential tool to prevent large-scale disasters [3]. Uncrewed aerial vehicle (UAV)-based detection is a promising technology with the potential to enable rapid deployment of surveys over large areas with minimal infrastructure. In remote areas, however, the UAVs are limited to onboard computing for detection due to the lack of high-bandwidth mobile networks.
@@
In this work, we propose using segmentation models capable of accurate real-time detection of wildfires with purely UAV-onboard computing resources. We leveraged bounding box-guided pseudo-label supervision from large foundation models to train the specialised models as simplified in Figure 1. We introduce two alternative approaches for supervising the final inference model using either a bounding box-guided zero-shot foundation model teacher or a larger model trained using state-of-the-art bounding box supervision methods. The methods were evaluated on a diverse human-annotated test set which showed their benefits compared to direct bounding box supervision. Real-world demonstrations and tests on additional data and hardware settings indicated the methods’ real-world applicability. The results suggest that largely used bounding box detection methods could be turned into segmentation at minimal costs.
@@ 2 Related Work
The ideas of UAV-based wildfire monitoring, detection and localisation date back to the early 2000’s [5, 38, 6]. However, at the dates of these studies, the available computational resources and non-learning-based computer vision methods posed serious limitations. More recent studies have applied bounding box detection to the task of wildfire detection [73, 50, 64, 53] and some of the most used bounding box detection models include R-CNN [19], Faster R-CNN [48], YOLO [47], with its numerous variants, and FPN [31]. Notably, the single-stage object detection networks, including YOLO and its variants focus on the real-time performance of the task and at best can work with very limited computational resources.
@@
The bounding box detections, however, can not fully capture the shape of the detected object which, in the case of UAV-based wildfire detection, could affect the localisation accuracy by even thousands of meters depending on the flight altitude and camera angle. This is key to our study, as we wish to develop a system capable of accurately estimating the 3D location of the fire onboard the UAV. Segmentation combats this issue by capturing the shape of the detected object with pixel-level accuracy. Recent segmentation models, such as ERFNet [49], BiSeNet [71], and ENet [43] have been proposed specifically for real-time use and segmentation has also been applied to wildfire detection [51, 72, 61]. However, while the studies have focused on UAV-captured images, their inference times have not been estimated with UAV-carried computers. Also, the studies mostly focused on flame detection but wildfire smoke is most often visible far further than any flames. The lack of public data, especially such, that is annotated at the pixel level, has largely limited studies on smoke segmentation.
@@
Publicly available wildfire datasets include the FLAME dataset [52], its synthetic extension FLAME Diffuser [60] and the Corsican dataset [57], all of which focus on flames. The only publicly available smoke focused datasets that could be found at the date of this study were the Bounding Box Annotated Wildfire Smoke Dataset Versions 1.0 and 2.0 by AI For Mankind and HPWREN [18] and a small dataset by the Croatian Center for Wildfire Research [28] of 98 images with varying quality pixel-level annotations.
@@
To tackle the lack of pixel-level annotated wildfire smoke data, we explored weak supervision, which could enable training segmentation models using other than dense pixel-level annotations. Weakly supervised segmentation refers to training segmentation models using non-pixel-level labels or far fewer labels than the number of input pixels. This can be realised with image-level labels, scribbles, individual pixels, or bounding boxes, as in the case of this study. The earlier solutions for box-supervised training of semantic and instance segmentation have used iterative mask proposals [13] or a combination of classical (non-learning-based) and learning-based solutions [26]. Many of the recent studies on weakly supervised segmentation have concentrated on instance segmentation which refers to models which differentiate between object instances without any semantic or object class information. Some of the recent proposals include Mask Auto-Labeler (MAL) [29], BoxTeacher [9], Semantic-aware Instance Mask Generation [30], and BoxSnake [69]. For wildfire detection, weakly supervised methods have been proposed to generate synthetic data [42, 74], to produce flame and smoke segmentation from classification models [1], and to train smoke classification and flame segmentation models [62]. While some of those models have been suggested for UAV-collected images, they were not tested for real-time sensing with limited hardware.
@@
To obtain real-time capable segmentation models, we suggest using knowledge distillation. Knowledge distillation is a way to train effective specialised networks by minimising the difference between the outputs of a larger model, assumed capable of better performance, and those of a smaller, distilled, model [22]. Knowledge distillation has also been applied to semantic segmentation model training for different applications in various forms with positive results [34, 44, 68, 67]. Knowledge distillation has also been proposed to improve wildfire detection [54, 65] but only for bounding box detection models.
@@
The recent emergence of prompt-guided segmentation [37] and specific models for the task, such as Segment Anything Model (SAM) [27] and High-Quality SAM (HQ-SAM) [25], provide new tools for generating pixel-level pseudo-labels from weak labels. The use of combined foundation model and weak label supervision has been shown successful with image level supervision [70, 55, 7, 8] alongside Class Activation Maps (CAMs) [75] or language understanding from language-image foundation models, such as Grounded-DINO [33] or CLIP (Contrastive Language-Image Pre-training) [45]. Foundation models have been applied to synthetic annotated wildfire flame image generation [60] but distilling the information to real-time models has not been studied nor the applicability to UAV-based systems in general. However, the use of combined bounding box and foundation model supervision for training use case-specific models has been proposed for planetary geological mapping [23], cell nuclei segmentation [11], and remote sensing [59]. For cell segmentation of histopathological imagery, such a model has been compared to methods that also use SAM at test time [58]. These studies suggest that such method could also be applied to train real-time wildfire smoke detection.
@@ 3 Methods
In this study, we address the gaps in the conducted literature by analysing the performance of PIDNet [66] models trained for the segmentation of wildfire smoke using labels obtained from a larger teacher model. We evaluated different teacher models for training the inference models by comparing the resulting accuracy of the final PIDNets. In the proposed framework, the teacher model is a Segment Anything Model guided by the bounding box labels to generate pseudo-labels for supervising the inference model. We compared this method to an alternative, where the teacher model is a larger binary segmentation model trained using state-of-the-art bounding box supervision methods.
We chose PIDNet as the inference architecture due to its capability of real-time performance even with limited computational resources. PIDNet is a three-branch model where each branch has their own focus: local semantic features, global semantic features, and boundary detection. The predictions are combined for the final inference segmentation output. The training method can be formulated as a knowledge distillation scheme with a teacher and a student model, shown in Figure 3.
@@ 3.1 Data
where $N$ denotes the total number of samples and $A_{n}$ and $B_{n}$ denote the predicted and the label masks of the individual samples respectively. In addition, the accuracy, precision, recall, and $F_{1}$ scores are reported in the main model comparison and their respective equations are shown in the Appendix Eqs. 3, 4, 5 and 6.
@@ 3.3 Inference model training
We used AdamW [36] with a starting learning rate of 1e-3 and weight decay coefficient of 1e-2 to optimise the inference models. The models were trained for 50 epochs and the validation loss and mean intersection over union was computed on each epoch, from which the best-performing model was chosen for evaluation on the test set and real-world inference based on the validation set mIoU. Minibatches of 16 images were used for the training and the models were pretrained on ImageNet [14].
@@
where loss functions $l_{0}$ and $l_{1}$ are binary cross entropy losses from the two alternative segmentation outputs of the PIDNet. $l_{2}$ is a weighted boundary-awareness cross entropy loss [56] and $l_{3}$ is a BAS-Loss that is a binary cross entropy loss computed only at the pixels where the boundary output of the model exceeds the threshold $t$. For the weighting of the losses we used the parameters $\lambda_{0}=0.4$, $\lambda_{1}=20$, $\lambda_{2}=1$, $\lambda_{3}=1$ and $t=0.8$ as set empirically in the study introducing the PIDNet model.
@@ 4 Results
Out of the PIDNet variants, PIDNet-M was found to be the best for the task regarding the test set metrics regardless of the teacher model. However, for practice, the PIDNet-S was chosen because of the far superior inference speed on the NVIDIA Jetson Orin NX, shown in Table 3 and the more extensive model speed benchmark with hardware setting variation in the Appendix Table 14. We expect the consistently worse performance of the PIDNet-L model to be caused by overfitting to noisy pseudo-labels, caused by the teacher models poorly capturing the human perception of the smoke in some images. Still, it is interesting how this behaviour was reflected in all teacher variants.
@@ 5 Conclusions
The proposed method also heavily relies on the quality of the SAM-generated pseudo-labels, which as indicated by the results, were not perfect. Thus, it would be of high value to further explore how the pseudo-label masks could be improved. Some simple strategies, such as adding variation to the prompts by random shifting of the manually generated bounding boxes or by assigning randomly chosen points inside the box as prompts could prove beneficial. The variationally generated masks could be used either as label augmentations or to produce more consistent pseudo-labels. Another possibility would be to fine-tune SAM specifically for the smoke segmentation task.
Acknowledgments
This research was funded by the Academy of Finland within project Fireman (decision no. 346710, 348009) and ML4DRONE (decision no. 357380). The FireMan project is funded under the EU’s Recovery and Resilience Facility that promotes the green and digital transitions through research. This study has been performed with affiliation to the Academy of Finland Flagship Forest–Human–Machine Interplay—Building Resilience, Redefining Value Networks and Enabling Meaningful Experiences (UNITE) (decision no. 357908).